OpenStudy (anonymous):

(-2,-2), m=-5 a.y=-5x-8 b.y=5x-12 c.y=-5x-12 d.y=-5x+12

OpenStudy (anonymous):

write an equation of the line that passes through each point, with the given slope

OpenStudy (anonymous):

equation of the line that passes through (x1,y1), with the given slope m \[\huge y-y1= m(x-x1)\]

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

now what is your choice?

OpenStudy (anonymous):

I figured it was y=-5x+12

OpenStudy (anonymous):

y-(-2)=-5(x-(-2))=-5(x+2)=-5x-10

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

thats not even an option though :P

OpenStudy (anonymous):

y+2=-5x-10 y=-5x-12